Given the function f is defined by f(x)= ax^2 + bx + c Given that f(0) = 6, f(-1)= 15, and f(1) = 1, find the values of a, b, and c.
Answer by shymeerab(21)

f(x)=ax^2+bx+c
for x=0
f(0)=c as given f(0)=6 so c=6
for x=-1
f(-1)=a-b+6 as given f(-1)=15 so a-b+6=15; a-b=9 _____e
for x=1
f(1)=a+b+6 as given f(1)=1 so a+b+6=1; a+b=-5 _____f
by adding eq e & f we have a=2
and put value of 'a' in eq f we have b=-7 so we have
a=2;b=-7;c=6